# Axis Tour: Word Tour Determines the Order of Axes in ICA-transformed Embeddings

### v1

@@ Abstract
Word embedding is one of the most important components in natural language processing, but interpreting high-dimensional embeddings remains a challenging problem. To address this problem, Independent Component Analysis (ICA) is identified as an effective solution. ICA-transformed word embeddings reveal interpretable semantic axes; however, the order of these axes are arbitrary. In this study, we focus on this property and propose a novel method, Axis Tour, which optimizes the order of the axes. Inspired by Word Tour, a one-dimensional word embedding method, we aim to improve the clarity of the word embedding space by maximizing the semantic continuity of the axes. Furthermore, we show through experiments on downstream tasks that Axis Tour constructs better low-dimensional embeddings compared to both PCA and ICA.
@@ 1 Introduction
Figure 1: Scatterplots of normalized ICA-transformed word embeddings whose axes are ordered by Axis Tour and Skewness Sort. In the upper part, Axis Tour is applied to 300-dimensional GloVe, with nine consecutive axes arranged counterclockwise. In the lower part, these nine axes are arranged clockwise in descending order of skewness. The embeddings are projected onto two dimensions along these axes. The top five embeddings on each axis are labeled by their words. Each word is assigned the color of the axis with the highest value. In both cases, words that cross the horizontal axes are not displayed.
Word embedding is one of the most important components in natural language processing, but interpreting high-dimensional embeddings remains a challenging problem. To address this problem, Independent Component Analysis (ICA) Hyvärinen and Oja (2000) is identified as an effective solution Mareček et al. (2020); Musil and Marecek (2022); Yamagiwa et al. (2023). ICA-transformed word embeddings reveal interpretable semantic axes; however, the order of these axes are arbitrary Hyvärinen et al. (2001b).
Inspired by a one-dimensional word embeddings, Word Tour Sato (2022), which leverages the Traveling Salesman Problem (TSP), we aim to improve the clarity of the word embedding space by maximizing the semantic continuity of the axes.
Fig. 1 shows two sets of two-dimensional projections of word embeddings: one set is ordered by Axis Tour and the other is sorted in descending order of skewness (Skewness Sort). In Axis Tour, the most significant words of the axes are far from the center, with the meanings of the axes changing continuously. Conversely, in Skewness Sort, the most significant words are nearer to the center, and axes with different meanings are placed adjacently. In fact, the average distance from the origin to these significant words in Fig. 1 is $0.76$ in Axis Tour, as opposed to $0.65$ in Skewness Sort.
We also assume that the consecutive axes in the Axis Tour embeddings can be considered as a subspace whose axes have similar meanings. Based on this, we project each subspace onto a single dimension to achieve dimensionality reduction. We show through experiments that Axis Tour yields better low-dimensional embeddings when compared to both PCA and ICA.
@@ 2 Related work
There are studies that transform embeddings by rotation Park et al. (2017) from Factor Analysis Crawford and Ferguson (1970); Browne (2001) or Principal Component Analysis (PCA) Musil (2019). Independent Component Analysis (ICA)  Hyvärinen and Oja (2000) has gained attention for its ability to reveal interpretable semantic axes in the transformed embeddings Mareček et al. (2020); Musil and Marecek (2022); Yamagiwa et al. (2023).
Relevant for our study is Topographic ICA (TICA) Kohonen (2001); Hyvärinen et al. (2001a). TICA relaxes the assumption of statistical independence and assumes energy correlations between adjacent axes. TICA then estimates the order of the axes. Unlike TICA, Axis Tour applies to ordinary ICA-transformed embeddings and uses the embeddings themselves to measure axis similarity.
@@ 3 Background
The pre-trained word embedding matrix is given by $\mathbf{X}=[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}]^{\top}\in\mathbb{R}^{n\times d}$, where $\mathbf{X}$ is centered (i.e., the mean of each column is zero). Here, $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents the word embedding of the $i$-th word.
@@ 3.1 ICA-transformed word embeddings
ICA Hyvärinen and Oja (2000) finds the transformation matrix $\mathbf{B}\in\mathbb{R}^{d\times d}$ such that the columns of the matrix $\mathbf{S}\in\mathbb{R}^{n\times d}$, represented by the following equation, are as statistically independent as possible:
@@
where $\mathbf{S}$ is assumed to be whitened (i.e., the variance of each column is $1$). The columns of $\mathbf{S}$ are called independent components [^1] Unless otherwise noted, flip the sign of each axis as needed so that the skewness is positive..
While the matrix $\mathbf{S}$ has interpretable semantic axes Mareček et al. (2020); Musil and Marecek (2022); Yamagiwa et al. (2023), the order of these axes are arbitrary Hyvärinen et al. (2001b).
@@ 3.2 Word Tour
$\underset{\sigma\in\mathcal{P}([n])}{\operatorname{min}}\,\|\mathbf{x}_{\sigma_{1}}-\mathbf{x}_{\sigma_{n}}\|+\sum_{i=1}^{n-1}\|\mathbf{x}_{\sigma_{i}}-\mathbf{x}_{\sigma_{i+1}}\|.$ (2)
@@ 4.1 Definition of axis embedding
We define axis embedding for use in Word Tour. The embedding represents the axis of the ICA-transformed embeddings $\mathbf{S}$.
In preparation, we define the matrix $\hat{\mathbf{S}}\in\mathbb{R}^{n\times d}$ as the normalization of the embeddings $\mathbf{S}$ with row vectors $\hat{\mathbf{s}}_{i}=\mathbf{s}_{i}/\|\mathbf{s}_{i}\|$. Here, the $i$-th word embedding of $\mathbf{S}$ and $\hat{\mathbf{S}}$ are denoted by $\mathbf{s}_{i},\hat{\mathbf{s}}_{i}\in\mathbb{R}^{d}$, respectively. We compare the elements of the $\ell$-th axis of $\hat{\mathbf{S}}$ and denote the index set of words corresponding to the top $k$ elements as $\text{Top}_{k}^{\ell}$.
We then define the $\ell$-th axis embedding $\mathbf{v}_{\ell}$ for $\mathbf{S}$ as follows:
@@
As we saw in Fig. 1, since the meaning of an axis can be interpreted from the top words, $\mathbf{v}_{\ell}$ can be considered as the embedding that represents the meaning of the $\ell$-th axis of $\mathbf{S}$.
@@ 4.2 Determining the order of axes
The sum of cosine similarities between adjacent axis embeddings can be considered as a measure of the semantic continuity of the axes. Thus, Axis Tour determines the order of the axes by maximizing this measure.
@@ 4.3 Dimensionality reduction
First, we consider dimensionality reduction for $I_{r}$. To do this, we define a vector $\mathbf{f}_{r}=(f_{r}^{(\ell)})_{\ell=1}^{d}\in\mathbb{R}_{\geq 0}^{d}$ of length $1$ for each $I_{r}$ as follows:
@@
where $\alpha\in\mathbb{R}_{\geq 0}$. Then $\mathbf{T}\mathbf{f}_{r}\in\mathbb{R}^{n}$ can be considered as a projection of the subspace spanned by the axes of $\mathbf{T}$ corresponding to $I_{r}$ onto a one-dimensional space. Fig. 5 in Appendix A shows the projection for three consecutive axes.
Next, we define the matrix $\mathbf{F}:=[\mathbf{f}_{1},\cdots,\mathbf{f}_{p}]\in\mathbb{R}^{d\times p}$. Then $\mathbf{T}\mathbf{F}\in\mathbb{R}^{n\times p}$ represents the concatenated projections, that is, a dimensionality reduction of the $d$-dimensional embeddings $\mathbf{T}$ to $p$ dimensions. For more details, refer to Appendix A.
@@ 5 Experiments
Table 1: Semantic continuity of axes via Axis Tour for normalized ICA-transformed embeddings. We apply Axis Tour to 300-dimensional GloVe and extract nine consecutive axes, then display the top five words for each axis.
Figure 2: Histogram of the cosine similarity between consecutive axis embeddings $\cos(\mathbf{v}_{i},\mathbf{v}_{i+1})$. As an additional baseline, we sampled 300 random words, ordered them, and showed the histogram of adjacent cosine similarities. The dashed lines indicates the average similarity for each method. The distribution for Axis Tour shifts towards a more positive mean, distancing itself from 0. Conversely, other distributions roughly follow the normal distribution with means that are close to 0.
[…]
Figure 3: Relationship between the skewness and the average of consecutive two cosines (i.e., $(\cos(\mathbf{v}_{i-1},\mathbf{v}_{i})+\cos(\mathbf{v}_{i},\mathbf{v}_{i+1}))/2$) for (a) Axis Tour and (b) Skewness Sort. The left shows the skewness and the average of two cosines for both $y$-axes, and the right shows the scatter plot of these values. Spearman’s rank correlation is 0.43 for Axis Tour, while it is 0.04 for Skewness Sort.
Figure 4: Dimensionality reduction performance for embeddings. Each value represents the average of 30 analogy tasks, 8 word similarity tasks, or 6 categorization tasks. For detailed experimental results, refer to Appendix B.
Similar to the Word Tour experiments, we used 300-dimensional GloVe  Pennington et al. (2014) with $n=40{,}000$, and the LKH solver [^5] The LKH solver is an implementation of Lin-Kernighan algorithm Lin and Kernighan (1973); Helsgaun (2000). Helsgaun (2018) as the TSP solver.
For ICA, we employed FastICA Hyvärinen (1999) from scikit-learn Pedregosa et al. (2011), setting iterations to $10{,}000$ and tolerance to $10^{-10}$, consistent with the setting in Yamagiwa et al. (2023). We computed the axis embeddings $\mathbf{v}_{\ell}$ in (3) with $k=100$. For baselines, we used whitened PCA-transformed embeddings [^6] Whitened ICA-transformed embeddings are obtained by applying an orthogonal matrix to these embeddings., along with two types of whitened ICA-transformed embeddings: Random Order, which randomizes the axes’ order in $\mathbf{S}$ [^7] In the embeddings matrix, we do not flip the sign of each axis to ensure positive skewness., and Skewness Sort, which arranges the axes of $\mathbf{S}$ in descending order of skewness. See Appendix C for additional experiments.
@@ 5.1 Qualitative observation
Table 1 shows examples of consecutive axes extracted using Axis Tour. These examples demonstrate the semantic similarity between adjacent axes and illustrate how the meanings of the axes change continuously. For example, in top row, the axis meaning transition from Eastern Europe to Germany and France, followed by Canada (which shares a connection with France) then to Australia (English-speaking regions), the regions in England, and finally to soccer (a popular sport in England), showcasing geographic and cultural continuity.
@@ 5.2 Cosine similarity between adjacent axis embeddings
Fig. 2 shows the histograms of $\cos(\mathbf{v}_{i},\mathbf{v}_{i+1})$ for Axis Tour and the baselines. In Axis Tour, the values of $\cos(\mathbf{v}_{i},\mathbf{v}_{i+1})$ are consistently higher, while this trend is not observed in the other baselines. This result is consistent with the formulation in (4).
For Axis Tour and Skewness Sort, Fig. 3 illustrates the relationship between the skewness and the average of the two adjacent cosine similarities [^8] Cosine similarity is not defined for the $i$-th axis itself.. In Fig. 3 for Axis Tour, it is interesting to note that these two values are correlated. However, no such trend is observed in Fig. 3 for Skewness Sort. This indicates that in Axis Tour, the axes with higher skewness are more likely to be ordered adjacent to the axes with higher similarity.
@@ 5.3 Dimensionality reduction: analogy, word similarity, and categorization tasks
[…]
Fig. 4 shows that the dimensionality reduction for Axis Tour is better than or equal to the baselines for most dimensions in each task. This result suggests that the dimensionality reduction for Axis Tour efficiently merges the axes with similar meanings. For more details, refer to Appendix B.
@@ 6 Conclusion
In this study, we propose a novel method, Axis Tour, which optimizes the order of axes in ICA-transformed word embeddings. We focus on the fact that the word embeddings reveal interpretable semantic axes while the order of these axes are arbitrary. Axis Tour aims to improve the clarity of the word embedding space by maximizing the semantic continuity of the axes. Furthermore, we show through experiments on downstream tasks that Axis Tour constructs better low-dimensional embeddings compared to both PCA and ICA.
@@
      Dimension reduction is performed using the vector $\mathbf{f}_{r}$, but its definition (5) is empirical, and better vectors may be designed. In addition, nonlinear transformations beyond linear ones could be considered for dimension reduction. Details on the definition of $\mathbf{f}_{r}$ can be found in Appendix A.
@@
This study complies with the ACL Ethics Policy.

### v2

@@ Abstract
Word embedding is one of the most important components in natural language processing, but interpreting high-dimensional embeddings remains a challenging problem. To address this problem, Independent Component Analysis (ICA) is identified as an effective solution. ICA-transformed word embeddings reveal interpretable semantic axes; however, the order of these axes are arbitrary. In this study, we focus on this property and propose a novel method, Axis Tour, which optimizes the order of the axes. Inspired by Word Tour, a one-dimensional word embedding method, we aim to improve the clarity of the word embedding space by maximizing the semantic continuity of the axes. Furthermore, we show through experiments on downstream tasks that Axis Tour yields better or comparable low-dimensional embeddings compared to both PCA and ICA.
@@ 1 Introduction
Figure 1: Scatterplots of normalized ICA-transformed word embeddings whose axes are ordered by Axis Tour and Skewness Sort. In the upper part, Axis Tour is applied to 300-dimensional GloVe, with nine consecutive axes arranged counterclockwise. In the lower part, these nine axes are arranged clockwise in descending order of skewness. The embeddings are projected onto two dimensions along these axes. The top five embeddings on each axis are labeled by their words. Each word is assigned the color of the axis with the highest value. In both cases, words that cross the horizontal axes are not displayed. Refer to Appendix A for more details.
Embedding is an important tool in natural language processing, but interpreting high-dimensional embeddings is challenging. To address this, Independent Component Analysis (ICA) Hyvärinen and Oja (2000) is an effective solution Mareček et al. (2020); Musil and Mareček (2022); Yamagiwa et al. (2023). ICA-transformed embeddings reveal interpretable semantic axes; however, the order of these axes are arbitrary Hyvärinen et al. (2001b). In this study, inspired by a one-dimensional word embeddings, Word Tour Sato (2022), which leverages the Traveling Salesman Problem (TSP), we aim to improve the clarity of the word embedding space by maximizing the semantic continuity of the axes.
Fig. 1 shows two sets of two-dimensional projections of word embeddings: one is ordered by Axis Tour and the other is sorted in descending order of skewness (Skewness Sort). In Axis Tour, the top words of the axes are far from the center, with the meanings of the axes changing continuously. Conversely, in Skewness Sort, the top words are nearer to the center, and the axes with different meanings are placed adjacently. In fact, the average distance from the origin to the top words in Fig. 1 is $0.76$ in Axis Tour, as opposed to $0.61$ in Skewness Sort.
We also assume that the consecutive axes in the Axis Tour embeddings can be considered as a subspace whose axes have similar meanings. Based on this, we project each subspace onto a single dimension for dimensionality reduction. We show through experiments that Axis Tour yields better or comparable low-dimensional embeddings compared to both PCA and ICA.
@@ 2 Related work
Some studies transform embeddings by rotation Park et al. (2017) from Factor Analysis Crawford and Ferguson (1970); Browne (2001) or Principal Component Analysis (PCA) Musil (2019). Independent Component Analysis (ICA) Hyvärinen and Oja (2000) has gained attention for its ability to reveal interpretable semantic axes in the transformed embeddings Mareček et al. (2020); Musil and Mareček (2022); Yamagiwa et al. (2023).
Relevant for our study is Topographic ICA (TICA) Kohonen (2001); Hyvärinen et al. (2001a). TICA relaxes the assumption of statistical independence and assumes higher-order correlations between adjacent axes. TICA then estimates the order of the axes. Unlike TICA, Axis Tour applies to ordinary ICA-transformed embeddings and uses the embeddings themselves to measure axis similarity. For more details on TICA, refer to Appendix F.
@@ 3 Background
The pre-trained word embedding matrix is given by $\mathbf{X}=[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}]^{\top}\in\mathbb{R}^{n\times d}$, where $\mathbf{X}$ is centered (i.e., the mean of each column is zero). Here, $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents the word embedding of the $i$th word.
@@ 3.1 ICA-transformed word embeddings
ICA Hyvärinen and Oja (2000) finds the transformation matrix $\mathbf{B}\in\mathbb{R}^{d\times d}$ such that the columns of the matrix $\mathbf{S}\in\mathbb{R}^{n\times d}$, represented by the following equation, are as independent as possible:
@@
where $\mathbf{S}$ is whitened (i.e., the variance of the columns is $1$ and their correlation is 00). The columns of $\mathbf{S}$ are called independent components [^1] Flip axis signs as needed to make the skewness positive.. While $\mathbf{S}$ has interpretable semantic axes Mareček et al. (2020); Musil and Mareček (2022); Yamagiwa et al. (2023), the order of these axes are arbitrary Hyvärinen et al. (2001b).
@@ 3.2 Word Tour
$\underset{\tau\in\mathcal{P}([n])}{\operatorname{min}}\,\|\mathbf{x}_{\tau_{1}}-\mathbf{x}_{\tau_{n}}\|+\sum_{i=1}^{n-1}\|\mathbf{x}_{\tau_{i}}-\mathbf{x}_{\tau_{i+1}}\|.$ (2)
@@ 4.1 Definition of axis embedding
We define axis embedding for use in Word Tour. The embedding represents the meaning of the axis of the ICA-transformed embeddings $\mathbf{S}$.
In preparation, we define the matrix $\hat{\mathbf{S}}\in\mathbb{R}^{n\times d}$ as the normalization of the embeddings $\mathbf{S}$ with row vectors $\hat{\mathbf{s}}_{i}=\mathbf{s}_{i}/\|\mathbf{s}_{i}\|$. Here, the $i$th word embedding of $\mathbf{S}$ and $\hat{\mathbf{S}}$ are denoted by $\mathbf{s}_{i},\hat{\mathbf{s}}_{i}\in\mathbb{R}^{d}$, respectively. We compare the elements of the $\ell$th axis of $\hat{\mathbf{S}}$ and denote the index set of words corresponding to the top $k$ elements as $\text{Top}_{k}^{\ell}$. We then define the $\ell$th axis embedding $\mathbf{v}_{\ell}$ for $\mathbf{S}$ as follows:
@@
As we saw in Fig. 1, since the meaning of an axis can be interpreted from the top words, $\mathbf{v}_{\ell}$ can be considered as the embedding that represents the meaning of the $\ell$th axis of $\mathbf{S}$.
@@ 4.2 Determining the order of axes
The sum of cosine similarities between adjacent axis embeddings can be considered as a metric of the semantic continuity of the axes. Thus, Axis Tour determines the order of the axes by maximizing this metric.
@@ 4.3 Dimensionality reduction
First, we consider dimensionality reduction for $I_{r}$. To do this, we define a unit vector $\mathbf{f}_{r}:=(f_{r}^{(\ell)})_{\ell=1}^{d}\in\mathbb{R}_{\geq 0}^{d}$ for each $I_{r}$ as follows:
@@
where $\alpha\in\mathbb{R}_{\geq 0}$. Then $\mathbf{T}\mathbf{f}_{r}\in\mathbb{R}^{n}$ can be considered as a projection of the subspace spanned by the axes of $\mathbf{T}$ corresponding to $I_{r}$ onto a one-dimensional space. Fig. 6 in Appendix B shows the projection for three consecutive axes.
Next, we define the matrix $\mathbf{F}:=[\mathbf{f}_{1},\cdots,\mathbf{f}_{p}]\in\mathbb{R}^{d\times p}$. Then $\mathbf{T}\mathbf{F}\in\mathbb{R}^{n\times p}$ represents the concatenated projections, that is, a dimensionality reduction of the $d$-dimensional embeddings $\mathbf{T}$ to $p$ dimensions. For more details, refer to Appendix B.
@@ 5 Experiments
Table 1: Semantic continuity of axes by Axis Tour for normalized ICA-transformed embeddings. We apply Axis Tour to 300-dimensional GloVe and show the top five words for each axis. See Appendix E.1 for all axis results.
Figure 2: Histogram of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$. As an additional baseline, we sampled 300 random words from the Random Order embeddings, ordered them, and showed the histogram. The dashed lines are the average similarity for each method. The distribution for Axis Tour shifts towards a more positive mean. Conversely, others roughly follow the normal distribution with means that are close to 0. For more details, refer to Appendix C.
(a) Axis Tour
(b) Skewness Sort
Figure 3: Relationship between the skewness and the average of consecutive two cosines (i.e., $(\cos(\mathbf{v}_{\ell-1},\mathbf{v}_{\ell})+\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1}))/2$) for (a) Axis Tour and (b) Skewness Sort. The left shows the skewness and the average of two cosines for both $y$-axes, and the right shows the scatter plot of these values. Spearman’s rank correlation is 0.43 for Axis Tour, while it is 0.04 for Skewness Sort.
Figure 4: The performance of dimensionality reduction for embeddings. Each value represents the average of 30 analogy tasks, 8 word similarity tasks, or 6 categorization tasks. See Appendix D for detailed experimental results.
Similar to the Word Tour experiments, we used 300-dimensional GloVe  Pennington et al. (2014) with $n=400{,}000$, and the LKH solver [^5] The LKH solver is an implementation of Lin-Kernighan algorithm Lin and Kernighan (1973); Helsgaun (2000). Helsgaun (2018) as the TSP solver.
For ICA, we used FastICA Hyvärinen (1999) from scikit-learn Pedregosa et al. (2011), setting iterations to $10{,}000$ and tolerance to $10^{-10}$, consistent with Yamagiwa et al. (2023). We computed the axis embeddings [^6] See Appendix E.3 for a discussion of the choice of $k$. $\mathbf{v}_{\ell}$ in (3) with $k=100$. For baselines, we used whitened PCA-transformed embeddings [^7] Whitened ICA-transformed embeddings are obtained by applying an orthogonal matrix to these embeddings., along with two types of whitened ICA-transformed embeddings [^8] See Appendix F for comparisons of Axis Tour and TICA.: Random Order, which randomly flips the sign of the axes in $\mathbf{S}$ and randomizes the order of the axes, and Skewness Sort, which sorts the axes of $\mathbf{S}$ in descending order of skewness. See Appendix E for additional experiments, including those of other embeddings.
@@ 5.1 Qualitative observation
Table 1 shows the illustrative examples of the meaning of countries, law, and art by the consecutive axes of the Axis Tour embeddings. These examples illustrate how the meanings of the axes change continuously. For example, in the top row, the axis meaning changes from Eastern Europe to Germany and France, followed by Canada (which shares a connection with France), then to Australia (English-speaking regions), the regions in England, and finally to soccer (a popular sport in England), demonstrating geographic and cultural continuity.
@@ 5.2 Cosine similarity between adjacent axis embeddings
Fig. 2 shows the histograms of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$ for Axis Tour and the baselines. In Axis Tour, the values of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$ are consistently higher, while this trend is not observed in the other baselines. This result is consistent with the formulation in (4).
For Axis Tour and Skewness Sort, Fig. 3 illustrates the relationship between the skewness and the average of consecutive two cosine similarities [^9] Cosine similarity is not defined for the $i$th axis itself.. In Fig. 3(a) for Axis Tour, it is interesting to note that these two values are correlated. However, no such trend is observed in Fig. 3(b) for Skewness Sort. This indicates that in Axis Tour, the axes with higher skewness are more likely to be ordered adjacent to the axes with higher similarity.
@@ 5.3 Dimensionality reduction: analogy, word similarity, and categorization tasks
Using Word Embedding Benchmark Jastrzebski et al. (2017) [^10] https://github.com/kudkudak/word-embeddings-benchmarks., we evaluated the performance of dimensionality reduction in analogy, word similarity, and categorization tasks. PCA selects the axes in descending order of eigenvalue. Random Order and Skewness Sort select the axes sequentially from the first to last. Axis Tour adopts the dimensionality reduction [^11] Fig. 7 in Appendix E.2 shows the results for different $\alpha$. in Section 4.3 with $\alpha=1/3$.
Fig. 4 shows that the dimensionality reduction for Axis Tour is better than or comparable to the baselines for most dimensions in each task. This result suggests that the dimensionality reduction for Axis Tour efficiently merges the axes with similar meanings. For more details, refer to Appendix D.
@@ 6 Conclusion
In this study, we proposed a novel method, Axis Tour, which optimizes the order of axes in ICA-transformed word embeddings. We focus on the fact that the word embeddings reveal interpretable semantic axes while the order of these axes are arbitrary. Axis Tour aims to improve the clarity of the word embedding space by maximizing the semantic continuity of the axes. We also showed through experiments on downstream tasks that Axis Tour yields better or comparable low-dimensional embeddings compared to both PCA and ICA.
@@
      Dimension reduction is performed using the vector $\mathbf{f}_{r}$, but its definition (5) is empirical, and better vectors may be designed. In addition, nonlinear transformations beyond linear ones could be considered for dimension reduction. Details on the definition of $\mathbf{f}_{r}$ can be found in Appendix B.
@@
A potential risk of this method is that we interpret the meanings of the axes of the ICA-transformed embeddings by the top words of each axis. If the embeddings contain personal information, such as email addresses or phone numbers, and these are contained in the top words, this can be problematic. Therefore, in this study, URLs, email addresses, and phone numbers were anonymized to avoid revealing such information.